# Unlocking Unlabeled Data: Ensemble Learning with the Hui-Walter Paradigm for Performance Estimation in Online and Static Settings

### v1

@@ Abstract
In the realm of machine learning and statistical modeling, practitioners often work under the assumption of accessible, static, labeled data for evaluation and training. However, this assumption often deviates from reality where data may be private, encrypted, difficult-to-measure, or unlabeled. In this paper, we bridge this gap by adapting the Hui-Walter paradigm, a method traditionally applied in epidemiology and medicine, to the field of machine learning. This approach enables us to estimate key performance metrics such as false positive rate, false negative rate, and priors in scenarios where no ground truth is available. We further extend this paradigm for handling online data, opening up new possibilities for dynamic data environments. Our methodology involves partitioning data into latent classes to simulate multiple data populations (if natural populations are unavailable) and independently training models to replicate multiple tests. By cross-tabulating binary outcomes across ensemble categorizers and multiple populations, we are able to estimate unknown parameters through Gibbs sampling, eliminating the need for ground-truth or labeled data. This paper showcases the potential of our methodology to transform machine learning practices by allowing for accurate model assessment under dynamic and uncertain data conditions.
@@ 1 Introduction
It is well known that data volume is growing rapidly. Most statistical, machine learning (ML), and deep learning techniques have developed when data were more scarce, and applications make increasingly unverified predictions based on data that is the same or similar to training data. Traditional statistical textbooks go to great lengths to emphasize the difference between interpolation and extrapolation, while warn of the dangers of the latter. Much of these data are delivered online (or streamed) and unlabeled. Researchers attempting to move machine learning online still assume that the data are mostly labeled (Gomes et al., 2019). Additionally, performance evaluation is a necessary and often daunting task when using a ML system in the production environment. Performance is often difficult to measure due to insufficient quantities and the variety of ground truth data labeled. Furthermore, data can be confidential, encrypted, or impossible (including costly) to measure, and the data landscape for online or streaming algorithms can constantly change. The actual distribution of the population is often unknown. Thus, it is impossible to relatively easily compare the output distribution of ML systems with the distribution of the population. Fortunately, the problems described are not limited to machine learning.
@@
Epidemiology is another discipline that relies on accurate efficacy measurements with insufficient ground truth and population data. Disease testing is the most recognizable example of epidemiology. Epidemiologists distinguish between efficacy and effectiveness when talking about moving from a laboratory environment where controlled experiments are possible (efficacy), whereby the test is deployed to test-recipient populations (effectiveness) (Singal et al., 2014). By analogy, a laboratory-developed Covid test measures efficacy by false-negative and false-positive rates in a double-blind controlled trial. The effectiveness is then measured by deploying the test to a remote population. In an applied data science environment, this mimics the applied data scientist who has trained a model in a notebook with labeled data and then deployed the model to production. Thus, the method we propose borrows concepts from how epidemiologists measure efficacy or effectiveness and apply them to measuring the performance of ML systems in production environments where there is no gold standard or ground truth. Our approach can be applied to any field in which unsupervised learning is ubiquitous and is generally likely to be of interest to anyone working in an applied data science environment. Alternatives to our approach are not common. Based on our research into this area, the most similar approach is analyzing agreement between multiple classifiers to arrive at an estimate of performance; we provide comparisons of our approach to this strategy.
@@
We use the Hui-Walter paradigm, which uses Bayesian methods to estimate the prior probability (prevalence) and false positive and false negative rates in ensemble models trained to detect the results of running in production or online (Hui & Walter, 1980). This estimate uses binary results of diagnostic tests, i.e. Covid positive or negative, for multiple tests and multiple populations in a table. We then performed Gibbs sampling to estimate unknown parameters, priors, false positive, and false negative rates without relying on ground truth data. We can make these estimates because more than one population provides us with enough equations and unknowns to solve the parameters of interest. We expand this idea by introducing a closed form solution in which two populations naturally occur, i.e. multiple data centers or regions, and use it to extend this framework to the online environment. When only one population is available, we use latent structure models for both online and offline methods. Many of the applications use bagging techniques or other ensemble methods in many settings. These ensemble methods naturally extend to our methodology. In our experiments, we divided the Wisconsin breast cancer and Adult data sets into latent classes for demonstration purposes, to simulate scenarios with multiple population datasets, and train independent models for simulation of multiple tests.
2 Proposed Method
@@
The Hui-Walter paradigm is a technique epidemiologists use to measure effectiveness without establishing the fundamental truth. The paradigm is based on the comparison of multiple tests and their results applied to more than one logical subset of the population (Hui & Walter, 1980). When evaluating ML systems, tests are analogous to binary categorizers, results are predictions (i.e. a $1$ or 00 binary classification corresponds to a positive or negative test result), and population subsets are groups of observations that are considered to be in distinct subpopulations. To obtain different subpopulations when only one exists, which may be the case when evaluating some ML systems, we propose using latent structural analysis, consisting of latent class analysis (LCA) or latent profile analysis (LPA), to split the data into two populations. Of course, some ML systems may already use multiple distinct subpopulations; for example, a practitioner has different data centers or databases to choose from. The framework we propose is two-fold. For machine learning practitioners working in an industry where multiple populations are often available for selection, we may consider the Hui-Walter paradigm applied to two or more models trained to detect results based on these data. If only one population is present, we use LPA or LCA to separate our data into multiple latent populations. Next, we give a closed form solution applying Hui-Walter to the data stream, measuring prior probabilities and false positive and false negative rates for unlabeled data streams. Our proposed solution dynamically assigns latent classes to streaming data and calculates unknown performance metrics for unlabeled data streams.
2.1 Review
2.1.1 Latent Class and Latent Profile Analysis
Latent structure analysis is a statistical method used to identify patterns or structures that underlie data. Latent class analysis is commonly used in psychology, sociology, and public health to identify subgroups or classes in a population based on their responses to a set of categorical variables. Similarly, researchers use latent profile analysis for continuous variables, and both models are generalized under the latent structure analysis moniker. They are also valuable tools for data scientists interested in understanding the underlying structure of a population and identifying subpopulations based on categorical, ordinal, or continuous variables Goodman (1974). In the analysis of latent profiles, we assume that each latent class follows a normal distribution of features. Note that the normality assumption only applies to these subpopulations and not to their combined distribution, which we model as a Gaussian mixture model. Each latent $(1,\dots,K)$ class has a density component in the mixed model.
2.1.2 Maximum Likelihood Estimate for Product Multinomial Distributions
Maximum likelihood estimation (MLE) is a method for estimating the parameters of a statistical model given a set of observations. One common application of MLE is to estimate the parameters of the multinomial distribution (Poleto et al., 2014).
The multinomial distribution is a generalization of the binomial distribution that allows for more than two possible outcomes. Given a set of $N$ independent trials, the multinomial distribution describes the probability of observing a specific combination of counts. The probability mass function of the multinomial distribution is given by $P(x_{1},x_{2},\dots,x_{k}|n,p_{1},p_{2},\dots,p_{k})=\frac{n!}{x_{1}!x_{2}!\dots x_{k}!}p_{1}^{x_{1}}p_{2}^{x_{2}}\dots p_{k}^{x_{k}}$, where $x_{i}$ is the number of times the $i$-th outcome was observed, $n$ is the total number of trials, and $p_{i}$ is the probability of the $i$-th outcome occurring.
To perform MLE on the multinomial distribution, we wish to find the values of $p_{1},p_{2},\dots,p_{k}$ that maximize the likelihood of the observed data. We can do this by taking the product of the probabilities of each outcome and maximizing this product for the parameters. This procedure is equivalent to maximizing the log-likelihood function $\mathcal{L}(\mathbf{p})=\sum_{i=1}^{k}x_{i}\log p_{i}$. The log-likelihood function is easier to work with because the product of many small numbers can be unstable, while the sum of the logs is more stable for calculation.
We can find the maximum likelihood estimates by setting the partial derivatives of the log-likelihood function to zero and solving for $p_{i}$. This reasoning leads to the following unbiased estimators $\hat{p}_{i}=\frac{x_{i}}{n}$.
The product-multinomial distribution is the distribution of multiple independent multinomial distributions. Extending the MLE to product-multinomial distributed data is as follows for $n$ independent multinomial distributions:
The likelihood function is given by $L(\bm{X})=\prod_{i=1}^{n}\prod_{j=1}^{k}p_{ij}^{x_{ij}}$ where $n$ is the number of populations, $k$ is the number of categories, $x_{ij}$ is the number of times the category $j$ was observed in the population $i$ and $\bm{X}$ is a vector of the parameters of the model.
The maximum likelihood estimate of the parameters, denoted by $\hat{\bm{X}}$, is found by maximizing the likelihood function of the parameters:
$\hat{\bm{X}}=\arg\max_{\bm{X}}L(\bm{X})$
We can estimate this using numerical optimization techniques.
The maximum likelihood estimate is not necessarily the best estimate of the parameters, and there are other methods of estimation, such as the method of moments and Bayesian estimation, that may be more appropriate in some cases. However, the maximum likelihood estimation is a widely used and well established method with several attractive properties, such as asymptotic efficiency (meaning that it becomes more accurate as the number of observations increases)(Hui & Walter, 1980).
In general, MLE is a valuable method for estimating the parameters of the product-multinomial distribution and has attractive theoretical properties. However, it is crucial to consider the method’s limitations, especially when working with small sample sizes.
2.1.3 Gibbs Sampling
Gibbs sampling, named after the physicist Josiah Willard Gibbs, is an analogy between the sampling algorithm and statistical physics. The algorithm was described by brothers Stuart and Donald Geman in 1984, some eight decades after the death of Gibbs, in their work on algorithmic image restoration (Geman & Geman, 1984). Gibbs sampling is a Markov chain Monte Carlo (MCMC) technique, typically the Metropolis-Hastings algorithm, for estimating the marginal distribution of variables in a multivariate distribution.
In Gibbs sampling, we start with an initial guess for the values of the variables and then iteratively sample each variable from its conditional distribution given the current values of the other variables. This process is repeated many times to obtain a sample from the joint distribution of the variables. Then we use the sample to estimate the marginal distribution of each variable or to compute other statistical quantities of interest.
Gibbs sampling has several advantages over other MCMC techniques. It is relatively simple to implement and easy to obtain good mixing (i.e., the convergence of the samples to the target distribution) in practice. It is also well suited to parallelization, which can speed up the sampling process.
However, Gibbs sampling can be inefficient when the variables are highly correlated or anticorrelated, as it can take many iterations to explore the full joint distribution. It is also sensitive to the choice of initial values and can get stuck in local modes for poorly chosen initial values.
Gibbs sampling finds applications in statistical physics, machine learning, and other fields where it is necessary to sample from multivariate distributions. It is beneficial for estimating the posterior distribution in Bayesian statistical models and is usually used with other MCMC techniques such as Metropolis-Hastings sampling.
Given a multivariate distribution, Gibbs sampling allows us to sample from a conditional distribution rather than to marginalize by integrating over a joint distribution; the joint distribution is often not known explicitly or difficult to sample from directly (Johnson et al., 2001). For purposes of our proposed approach in the next section, Gibbs sampling allows us to assume prior uncertainty for the prevalence (i.e. base rate) and classifier performance parameters (false positive rates, false negative rates) and represent them as independent distributions.
The algorithm is shown in Algorithm 1:
Algorithm 1 Gibbs Sampling
1:procedure GibbsSample($x_{1},\dots,x_{n},T$)
2:     Initialize $X^{(0)}=(x_{1},\dots,x_{n})$
3:     for $t=1,2,\dots,T$ do
4:         for $i=1,\dots,n$ do
5:              Sample $X^{(t)}_{i}$ from $p(X^{(t)}_{i}\mid X^{(t)}_{1},\dots,X^{(t)}_{i-1},X^{(t-1)}_{i+1},\dots,X^{(t-1)}_{n})$
6:         end for
7:     end for
8:     return samples $X^{(0)},\dots,X^{(T)}$
9:end procedure
This algorithm takes a set of variables $X=(x_{1},\dots,x_{n})$ and $T$ as input and returns $T$ samples from their joint distribution. It does this by iteratively sampling each variable from its conditional distribution given the current values of the other variables. The algorithm ends after a fixed number of iterations $T$.
2.2 Proposed
2.2.1 Hui-Walter
@@
The Hui-Walter paradigm is a framework from epidemiology for estimating the false positive rate, false negative rate, and prior probabilities across multiple populations, with more than one test. In the epidemiological context, the Hui-Walter paradigm was designed for disease tests, i.e. Covid tests, where the results are 00 or $1$ for does not-have-the-disease and has-the-disease, respectively. Since we are applying this paradigm to the machine learning context, we will discuss the method in terms of classifications from binary categorizers. The Hui-Walter paradigm suggests that we can form a $n\times m\times k$ contingency table for $n$ binary categorizers, $m$ populations, and $k$ classes. The goal of this method is to estimate a parameter vector $(\hat{\alpha},\hat{\beta},\hat{\theta})$ for each model and population, where $\alpha$ is the false positive rate or $1-\text{specificity}$, $\beta$ is the false negative rate or $1-\text{sensitivity}$, and $\theta$ is the prior probability of the positive class in the population. The Hui-Walter paradigm is valid for $m$ populations, $n$ classes, and $k$ categorizers where $m,n,k\in\mathbb{N}$. When $m=n=k=2$, we can show this result using maximum likelihood estimation Hui & Walter (1980).
@@
Now, the advantage of maximum likelihood for a single population is that $p_{1}$ is the probability that both tests produce a positive result. We solve this over-determined system by minimizing the likelihood function. In our simplified setup, we have a parameter vector $p=(p_{1},p_{2},p_{3},p_{4})$, which we wish to estimate. Because these probabilities in question are the outcomes of binary classifiers or diagnostic tests, we may express them in terms of the false positive and false negative rates of the models in question and the population’s prior probability. Let $\alpha_{1},\alpha_{2}$ be the false positive rates for $T_{1}$ and $T_{2}$, let $\beta_{1},\beta_{1}$ be the false negative rates for $T_{1}$ and $T_{2}$, and let $\theta$ be the prior probability for the positive class (base rate) of the population (or the disease in the epidemiological case).
@@
We need to consider two populations because the base rates of a single population are unknown. We assume that $\alpha_{i}$ and $\beta_{i}$ are the same in both populations for $i=1,2$, therefore we have six equations with six unknowns. In other words, $p_{i}$ where $i=1,2,\dots 8$, can be written algebraically as combinations of 6 variables. Therefore, we use two populations so that the above estimates are solvable (Johnson et al., 2001). At least two populations guarantee that we have as many equations as unknowns. For two populations, we have two independent multinomial distributions. Therefore, the likelihood function is the product of the likelihood functions (Hui & Walter, 1980). Now consider two populations, $P_{1}$ and $P_{2}$ with prior probabilities $\theta_{1}$ and $\theta_{2}$. In that case, our two-by-two-cell data becomes Figure 2.
@@
We want to estimate the parameter vector $\bm{X}=(\mathbf{\alpha},\mathbf{\beta},\mathbf{\theta})$. A real solution with coordinates that satisfy these constraints is a maximum likelihood estimate.
For our simplified case with two populations we get $\frac{\partial L}{\partial p_{i}}=\frac{x_{i}}{p_{i}}-\lambda=0,\frac{\partial L}{\partial p_{j}}=\frac{x_{j}}{p_{j}}-\mu=0$ for $i=1,\dots,4$, and for $j=5,\dots,8$. The maximum likelihood then is $\mathbf{p}=(\frac{p_{1}}{n_{1}},\dots,\frac{p_{4}}{n_{1}},\frac{p_{5}}{n_{2}},\dots,\frac{p_{8}}{n_{2}})$ from 1. This closed formula gives the maximum likelihood estimate for the product-multinomial distribution. This MLE extends naturally to setups with more than 2 classes.
@@
$\mathcal{I}(\bm{X})=\operatorname{\mathbb{E}}[\mathbf{Hess}(\mathbf{X}\otimes\mathbf{X})]$
In other words, when the MLE resides within the interior point of the feasible set, we can solve for the information matrix by taking the expectation value of the hessian of the outer product of our parameter vector, $\bm{X}$ (Blasques et al., 2018).
Inverting this matrix gives the variance and covariance matrix. Now that we have shown that there is a maximum likelihood estimate for the parameter vector $(\mathbf{\alpha},\mathbf{\beta},\mathbf{\theta})$ exists for the $2\times 2\times 2$ case, it is trivial to see that this generalizes to the $n\times m\times k$ case for arbitrary dimensions of the population outcome tensor. We can generally find a solution to the parameter vector using Bayesian methods such as Monte Carlo estimates or Gibbs sampling to estimate our parameter vector, $\bm{X}=(\mathbf{\alpha},\mathbf{\beta},\mathbf{\theta})$.
One of the assumptions of Hui-Walter is that the tests (or ML binary classifiers) be conditionally independent Hui & Walter (1980). We can test for conditional independence, which can easily be tested by performing a goodness-of-fit hypothesis test on the contingency table with a $G$-test (Hui & Zhou, 1998).
2.2.2 Hui-Walter Online
Online Machine Learning refers to training machine learning models on data that arrive continuously and in real time rather than using a static data set. In this setting, the model updates its parameters as new data arrive and makes predictions accordingly (Gomes et al., 2019). This approach is helpful in applications such as streaming data analysis, recommendation systems, and dynamic pricing.

### v2

@@ Abstract
In the industrial practice of machine learning and statistical modeling, practitioners often work under the assumption of accessible, static, labeled data for evaluation and training. However, this assumption often deviates from reality, where data may be private, encrypted, difficult-to-measure, or unlabeled. In this paper, we bridge this gap by adapting the Hui-Walter paradigm, a method traditionally applied in epidemiology and medicine, to the field of machine learning. This approach enables us to estimate key performance metrics such as false positive rate, false negative rate, and priors in scenarios where no ground truth is available. We further extend this paradigm for handling online data, opening up new possibilities for dynamic data environments. Our methodology involves partitioning data into latent classes to simulate multiple data populations (if natural populations are unavailable) and independently training models to replicate multiple tests. By cross-tabulating binary outcomes across multiple categorizers and multiple populations, we are able to estimate unknown parameters through Gibbs sampling, eliminating the need for ground-truth or labeled data. This paper showcases the potential of our methodology to transform machine learning practices by allowing for accurate model assessment under dynamic and uncertain data conditions.
keywords:
Machine Learning, Online Machine Learning, Bayesian Statistics, Data Science, Hui-Walter Paradigm, Unlabeled Data, Performance Metrics
@@ 1 Introduction
It is well known that data volume is growing rapidly. Most statistical, machine learning (ML), and deep learning techniques have developed when data were more scarce, and applications make increasingly unverified predictions based on data that is the same or similar to training data. Traditional statistical textbooks go to great lengths to emphasize the difference between interpolation and extrapolation, while warning of the dangers of the latter. Much of these data are delivered online (or streamed) and unlabeled. Researchers trying to move machine learning online still assume that the data are mostly labeled [17, 31]. Additionally, performance evaluation is a necessary and often daunting task when using a ML system in the production environment. Performance is often difficult to measure due to insufficient quantities and the variety of ground truth data labeled. Furthermore, data can be confidential, encrypted, or impossible (including costly) to measure, and the data landscape for online or streaming algorithms can constantly change. The actual distribution of the population is often unknown. Thus, it is impossible to relatively easily compare the output distribution of ML systems with the distribution of the population. Fortunately, the problems described are not limited to machine learning.
@@
Epidemiology is another discipline that relies on accurate efficacy measurements with insufficient ground truth and population data. Disease testing is the most recognizable example of epidemiology. Epidemiologists distinguish between efficacy and effectiveness when talking about moving from a laboratory environment where controlled experiments are possible (efficacy), whereby the test is deployed to test-recipient populations (effectiveness) [32]. By analogy, a laboratory-developed Covid test measures efficacy by false-negative and false-positive rates in a double-blind controlled trial. The effectiveness is then measured by deploying the test to a remote population. In an applied data science environment, this mimics the applied data scientist who trained a model in a notebook with labeled data and then deployed the model to production. Thus, the method we propose borrows concepts from how epidemiologists measure efficacy or effectiveness and apply them to measuring the performance of ML systems in production environments where there is no gold standard or ground truth. Our approach can be applied to any field in which unsupervised learning is ubiquitous and is generally likely to be of interest to anyone working in an applied data science environment. Alternatives to our approach are not common. Based on our research in this area, the most similar approach is to analyze the agreement between multiple classifiers to arrive at an estimate of performance; we provide comparisons of our approach to this strategy.
@@
We use the Hui-Walter paradigm, which uses Bayesian methods to estimate the prior probability (prevalence) and false positive and false negative rates in ensemble models trained to detect the results of running in production or online [19]. This estimate uses binary results of diagnostic tests, i.e. Covid positive or negative, for multiple tests and multiple populations in a table. We then performed Gibbs sampling [16, 15] to estimate unknown parameters, priors, false positive, and false negative rates without relying on ground truth data. We can make these estimates because more than one population provides us with enough equations and unknowns to solve the parameters of interest. We expand this idea by introducing a closed form solution in which two populations naturally occur, i.e. multiple data centers or regions, and use it to extend this framework to the online environment. When only one population is available, we use latent structure models for both online and offline methods. Many applications use bagging techniques or other ensemble methods in many settings. These ensemble methods naturally extend to our methodology. In our experiments, we divided the Wisconsin breast cancer and Adult data sets into latent classes for demonstration purposes, to simulate scenarios with multiple population data sets, and train independent models for simulation of multiple tests.
2 Related Work
The paper “Estimating Accuracy from Unlabeled Data: a Bayesian Approach” presents methods for estimating classifier accuracies using unlabeled data, focusing on agreement rate-based methods and Bayesian graphical models. Their method relies on agreement rates between classifiers, assuming independent errors, and operates within a static data environment. While effective for static data sets, their framework does not naturally extend to online or dynamic data streams. In contrast, Slote and Lee also assume independent classifiers but adapt the Hui-Walter paradigm from epidemiology, which allows for performance estimation across multiple populations or latent classes. Our work employs a complementary approach that estimates accuracy but focuses on multiple classifiers and multiple populations in an online and static setting, which could offer a practical variation for similar unsupervised learning tasks [28]. Due to the differences in approach, we employ the Rand index (RI) for comparison to our model.
3 Main results
@@
The Hui-Walter paradigm is a technique epidemiologists use to measure effectiveness without establishing the ground truth on new data. The paradigm is based on the comparison of multiple tests and their results applied to more than one logical subset of the population [19]. When evaluating ML systems, tests are analogous to binary categorizers, results are predictions (i.e. a $1$ or 00 binary classification corresponds to a positive or negative test result), and population subsets are groups of observations that are considered to be in distinct subpopulations. To obtain different subpopulations when only one exists, which may be the case when evaluating some ML systems, we propose using latent structural analysis, consisting of latent class analysis (LCA) or latent profile analysis (LPA), to split the data into two populations. Of course, some ML systems may already use multiple distinct subpopulations; for example, a practitioner has different data centers or databases to choose from. The framework we propose is two-fold. For machine learning practitioners working in an industry where multiple populations are often available for selection, we may consider the Hui-Walter paradigm applied to two or more models trained to detect results based on these data. If only one population is present, we use LPA or LCA to separate our data into multiple latent populations. Next, we give a closed-form solution applying Hui-Walter to the data stream, measuring prior probabilities and false positive and false negative rates for unlabeled data streams. Our proposed solution dynamically assigns latent classes to streaming data and calculates unknown performance metrics for unlabeled data streams.
4 Hui-Walter
@@
The Hui-Walter paradigm suggests that we can form a $n\times m\times k$ contingency table for $n$ binary categorizers, $m$ populations, and $k$ classes. The goal of this method is to estimate a parameter vector $(\hat{\alpha},\hat{\beta},\hat{\theta})$ for each model and population, where $\alpha$ is the false positive rate or $1-\text{specificity}$, $\beta$ is the false negative rate or $1-\text{sensitivity}$, and $\theta$ is the prior probability of the positive class in the population. The Hui-Walter paradigm is valid for $m$ populations, $n$ classes, and $k$ categorizers where $m,n,k\in\mathbb{N}$. When $m=n=k=2$, we can show this result using maximum likelihood estimation [19].
@@
Now, the advantage of maximum likelihood for a single population is that $p_{1}$ is the probability that both tests produce a positive result. We solve this overdetermined system by minimizing the likelihood function. In our simplified setup, we have a parameter vector $p=(p_{1},p_{2},p_{3},p_{4})$, which we wish to estimate. Because these probabilities in question are the outcomes of binary classifiers or diagnostic tests, we may express them in terms of the false positive and false negative rates of the models in question and the population’s prior probability. Let $\alpha_{1},\alpha_{2}$ be the false positive rates for $T_{1}$ and $T_{2}$, let $\beta_{1},\beta_{2}$ be the false negative rates for $T_{1}$ and $T_{2}$, and let $\theta$ be the prior probability for the positive class (base rate) of the population (or the disease in the epidemiological case).
@@
We need to consider two populations because the base rates of a single population are unknown. We assume that $\alpha_{i}$ and $\beta_{i}$ are the same in both populations for $i=1,2$, therefore we have six equations with six unknowns. In other words, $p_{i}$ where $i=1,2,\dots 8$, can be written algebraically as combinations of 6 variables. Therefore, we use two populations so that the above estimates are solvable [22]. At least two populations guarantee that we have as many equations as unknowns. For two populations, we have two independent multinomial distributions. Therefore, the likelihood function is the product of the likelihood functions [19]. Now consider two populations, $P_{1}$ and $P_{2}$ with prior probabilities $\theta_{1}$ and $\theta_{2}$. In that case, our two-by-two-cell data becomes Figure 2.
@@
We want to estimate the parameter vector $\boldsymbol{X}=(\mathbf{\alpha},\mathbf{\beta},\mathbf{\theta})$. A real solution with coordinates that satisfy these constraints is a maximum likelihood estimate.
For our simplified case with two populations we get
$\frac{\partial L}{\partial p_{i}}=\frac{x_{i}}{p_{i}}-\lambda=0,\frac{\partial L}{\partial p_{j}}=\frac{x_{j}}{p_{j}}-\mu=0$
for $i=1,\dots,4$, and for $j=5,\dots,8$. The maximum likelihood then is $\mathbf{p}=(\frac{p_{1}}{n_{1}},\dots,\frac{p_{4}}{n_{1}},\frac{p_{5}}{n_{2}},\dots,\frac{p_{8}}{n_{2}})$ from equation 1. This closed formula gives the maximum likelihood estimate for the product-multinomial distribution. This MLE extends naturally to setups with more than 2 classes.
@@
$\mathcal{I}(\boldsymbol{X})=\mathbb{E}\left[\operatorname{Hess}(\boldsymbol{X}\otimes\boldsymbol{X})\right]\text{.}$
In other words, when the MLE resides within the interior point of the feasible set, we can solve for the information matrix by taking the expectation value of the hessian of the outer product of our parameter vector, $\boldsymbol{X}$ [5].
Inverting this matrix gives the variance and covariance matrix. Now that we have shown that there is a maximum likelihood estimate for the parameter vector $(\mathbf{\alpha},\mathbf{\beta},\mathbf{\theta})$ exists for the $2\times 2\times 2$ case, it is trivial to see that this generalizes to the $n\times m\times k$ case for arbitrary dimensions of the population outcome tensor. We can generally find a solution to the parameter vector using Bayesian methods such as Monte Carlo estimates or Gibbs sampling to estimate our parameter vector, $\boldsymbol{X}=(\mathbf{\alpha},\mathbf{\beta},\mathbf{\theta})$.
One of the assumptions of Hui-Walter is that the tests (or ML binary classifiers) be conditionally independent [19]. We can test for conditional independence, which can easily be tested by performing a goodness-of-fit hypothesis test on the contingency table with a $G$-test [20].
5 Hui-Walter Online
Online Machine Learning refers to training machine learning models on data that arrive continuously and in real time rather than using a static data set. In this setting, the model updates its parameters as new data arrive and makes predictions accordingly [17]. This approach is helpful in applications such as streaming data analysis, recommendation systems, and dynamic pricing.